OpenStudy (anonymous):

jess had a balance of 1200 on a credit card with an 18.7% compound monthly.ABOUT HOW MUCH WILL HE SAVE IN ITEREST OVER THE COURSE OF A YEAR IF HE TRANSFER HIS BALANCE TO A CREDIT CARD WITH A 12.5% COMPOUND MONTHLY.

OpenStudy (hakunamatata34):

I can help

OpenStudy (anonymous):

ok,

OpenStudy (hakunamatata34):

ok when you divide 1200/18.7% you get $64.17 and when you divide 1200/ %12.5 you get $96.

OpenStudy (hakunamatata34):

am i wrong ?? Because that doesnt sound right to me??

OpenStudy (amistre64):

how do we calculate for interest to start with?

OpenStudy (hakunamatata34):

don't we divide 1200/18.7 or am I wrong??

OpenStudy (amistre64):

we are compounding monthly

OpenStudy (anonymous):

that's what I'm thinking

OpenStudy (amistre64):

I = P-P(1+r/12)^12

OpenStudy (hakunamatata34):

I'm in 9th grade I don't think i got to compound yet

OpenStudy (amistre64):

might have the terms backwards ... I1 = P(1+r1/12)^12 - P I2 = P(1+r2/12)^12 - P I1-I2 P(1+r1/12)^12 -P(1+r2/12)^12 P [(1+r1/12)^12 -(1+r2/12)^12]

OpenStudy (anonymous):

What is that?

OpenStudy (amistre64):

the interest calculations ...

OpenStudy (amistre64):

how do you calculate the interest compounding?

OpenStudy (anonymous):

can you show me the calculations because I don't know.

OpenStudy (amistre64):

the compound formula should be in your materials ... what is the formula for compounding interest?

OpenStudy (anonymous):

this so over my head lol